### v1

# A Classification Benchmark for Artificial Intelligence Detection of Laryngeal Cancer from Patient Speech

Abstract
Cases of laryngeal cancer are predicted to rise significantly in the coming years. Current diagnostic pathways cause many patients to be incorrectly referred to urgent suspected cancer pathways, putting undue stress on both patients and the medical system.
Artificial intelligence offers a promising solution by enabling non-invasive detection of laryngeal cancer from patient speech, which could help prioritise referrals more effectively and reduce inappropriate referrals of non-cancer patients. To realise this potential, open science is crucial. A major barrier in this field is the lack of open-source datasets and reproducible benchmarks, forcing researchers to start from scratch. Our work addresses this challenge by introducing a benchmark suite comprising 36 models trained and evaluated on open-source datasets. These models are accessible in a public repository, providing a foundation for future research. They evaluate three different algorithms and three audio feature sets, offering a comprehensive benchmarking framework. We propose standardised metrics and evaluation methodologies to ensure consistent and comparable results across future studies.
The presented models include both audio-only inputs and multimodal inputs that incorporate demographic and symptom data, enabling their application to datasets with diverse patient information. By providing these benchmarks, future researchers can evaluate their datasets, refine the models, and use them as a foundation for more advanced approaches. This work aims to provide a baseline for establishing reproducible benchmarks, enabling researchers to compare new methods against these standards and ultimately advancing the development of AI tools for detecting laryngeal cancer.
keywords:
Artificial Intelligence , Laryngeal Cancer , Benign Voice Disorders , Voice Classification
journal: arxiv
[…]
[label1]organization=School of Computer Science, University of Leeds,city=Leeds, postcode=LS2 9JT, country=UK
[…]
[label2]organization=Leeds Teaching Hospitals NHS Trust,city=Leeds, postcode=LS1 3EX, country=UK
[…]
[label3]organization=School of Mathematics, University of Leeds,city=Leeds, postcode=LS2 9JT, country=UK
1 Introduction
In 2022, over 188,000 people were diagnosed with laryngeal cancer worldwide. With cases predicted to rise by 66.2% by 2050, efficient diagnosis is necessary to reduce mortality rates (Bray et al., 2024). Early detection of laryngeal cancer improves a patient’s survival rates and increases treatment options, allowing for a better quality of life post-treatment (Cancer Research UK, 2019, 2021b). Current diagnostic procedures require invasive techniques, such as nasendoscopy or laryngoscopy, to obtain specimens for biopsy. Nasendoscopy, performed as an outpatient procedure, uses a small fibre-optic camera to view the larynx via the nose, while a direct laryngoscopy is performed under general anaesthetic in the operating theatre (NHS, 2017). These methods are uncomfortable, invasive and resource-intensive.
Artificial intelligence (AI) analysis of voice has been suggested as a non-invasive screening tool for the detection of laryngeal cancer, reducing the need for invasive and uncomfortable medical tests. Such a tool may be able to screen patients with concerns regarding their voice, prioritise those at the highest risk of a cancer diagnosis, expedite their specific care pathway, and increase the accessibility of diagnosis by reducing the need for expensive medical equipment as well as reducing patient stress and the load on the medical system.
Despite prior research on detecting laryngeal cancer from patient speech, progress in this field has been limited by the scarce availability of code, models and data which are rarely made available (Wang et al., 2024; Kim et al., 2024; Chen et al., 2023). Current approaches are ad-hoc with no standardisation in model assessment, making results incomparable and inconsistent and setting back advancements in this field (Paterson et al., 2024).
This work produces a suite of 36 models resulting from the combination of three algorithms, three audio feature sets, and four combinations of input variables ($3\times 3\times 4=36$). All 36 models perform the binary classification of benign and malignant voice pathologies. All pre-trained models and code used for training have been made publicly available. We also define and justify evaluation metrics, including classification performance, inference times, and fairness testing, to standardise assessment in this domain.
This work establishes a benchmark for future research, aiming to accelerate progress in developing generalizable and high-performing models for laryngeal cancer detection. Researchers can use this benchmark to evaluate new datasets, refine models, and overall improve progress in this area.
The rest of this work is structured as follows: Section 2 comprehensively describes the two datasets used in this work. Section 3 explains the methods used to develop the classifiers, including audio feature extraction and algorithm choice. This section also describes the evaluation methods used in this work and suggested for future work. Section 4 presents our results and discusses the impact of different algorithms, audio features, and demographic and symptom data incorporation. Section 5 concludes this work and discusses future work.
2 Evaluative Datasets
There are a limited number of publicly available datasets in this field. In our work, we use the only two datasets that, to the best of our knowledge, are currently available for public use: the Far Eastern Memorial Hospital voice dataset (FEMH) for model training and testing and the Saarbruecken Voice Database (SVD) for external validation. This section comprehensively describes and compares these two datasets.
2.1 Far Eastern Memorial Hospital voice dataset
[…]
Figure 1: The number of patients for each diagnosis split into benign and malignant.
We split this dataset using 33% of the data for testing (635 benign, 25 malignant) and the remaining data for training (1305 benign, 35 malignant). Table 2 shows a breakdown of the samples in the testing and training sets, including the number of male and female patients and the minimum, mean, and maximum age. All recordings were taken with a sample rate of 44,100 Hz.
2.2 Saarbruecken Voice Database
The Saarbruecken Voice Database (SVD) is an open-source dataset containing over 2000 participants and over 70 different pathologies (Manfred Pützer and William, J Barry, 2007). This dataset includes recordings of patients sustaining several vowels and saying a short sentence; in this work, we use the recordings of sustained /a/ at a natural pitch and volume for consistency with the audio in the FEMH dataset. This dataset also contains the patient’s age and gender.
An experienced clinician classified all of the pathologies in the dataset and identified eight malignant and pre-malignant conditions in the dataset. Table 1 shows these eight pathologies and the number of patients in each (total of 38 malignant patients). The remaining pathologies were used to form the benign pathology group, which contains 1301 patients. To ensure consistency with the data provided in the FEMH dataset, healthy participants and participants under the age of 18 were excluded in this work. These recordings were all taken between 1997 and 2004. All recordings were taken with a sample rate of 50,000 Hz.
Condition Condition Type Male Female Total
Vocal cord cancer Malignant 21 1 22
Hypopharyngeal tumor Malignant 6 0 6
Larynx tumor Malignant 4 1 5
Epiglottic cancer Malignant 0 1 1
Nesopharyngeal tumor Malignant 1 0 1
Carcinoma in situ Malignant 1 0 1
Dysplastic dysphonia Pre-malignant 1 0 1
Dysplastic larynx Pre-malignant 1 0 1
Table 1: The number of patients in the SVD per condition. This is also split into the number of male and female patients per condition.
Dataset Pathology Sex Count Min Age Mean Age Max Age
FEMH Train Benign Female 808 20 47 93
Male 497 20 53 97
Malignant Female 3 68 73 81
Male 32 31 64 88
FEMH Test Benign Female 379 20 45 88
Male 256 20 52 89
Malignant Female 3 50 58 63
Male 22 45 64 91
SVD Test Benign Female 716 18 49 94
Male 579 18 54 89
Malignant Female 3 51 54 58
Male 35 38 60 75
Table 2: Demographics in the training and test sets.
2.3 Dataset comparison
It’s important to compare the two datasets used in this work to understand how potential differences may impact the classifier performance.
We first compare the age and gender of the patients in each dataset. Figure 2(a) shows the distribution of ages in the datasets. It can be seen that the patients within the set of malignant samples are older than those in the benign group. Using a Mann-Whitney U test, the age distributions in the benign and malignant classes are compared between the two datasets (0.05 significance threshold). For the benign group, there is a significant difference between the age distributions (p=5.88e-6). For the malignant group, no significant difference is found (p=0.093).
Figure 2(b) shows the percentage of male and female patients within each group for each dataset. There are many more male than female patients in the malignant group. This is to be expected since laryngeal cancer affects many more men than women on average (Cancer Research UK, 2021a). Using a Fisher-Exact test, we compare the proportions of male and female patients in the benign and malignant classes between the two datasets (0.05 significance threshold). For the benign group, there is a significant difference between the proportion of male and female patients (p=9.06e-4). For the malignant group, no significant difference is found (p=1.0).
(a) The distribution of ages in the different datasets for the benign and malignant samples.
(b) The percentages of male and female samples in the different datasets for the benign and malignant samples.
Figure 2: A comparison of patient demographics within the FEMH train and set set alongside the SVD extrenal test set.
The two datasets contain a similar frequency of malignant patients, with malignant patients making up 3.0% of the FEMH dataset and 2.8% of SVD.
The length of audio recordings also varies significantly between the two datasets. The recordings within the FEMH dataset are all exactly one, one and a half, two, or three seconds long, with the majority being three seconds. The SVD recordings, however, vary more significantly between 0.4 and 2.6 seconds long, with a mean duration of 1.3 seconds. Differences in recording length could impact classification performance. Shorter recordings in the SVD may contain less information, which could lead to lower classification accuracy compared to the longer recordings in the FEMH dataset. Using a Mann-Whitney U test, the distribution of recording lengths between the benign and malignant classes are compared for each of the two datasets (0.05 significance threshold). There is no significant difference in the recording lengths between the malignant and benign patients in either dataset (p=0.962 for FEMH, p=0.761 for SVD).
3 Methods
3.1 Audio Feature Sets
Our benchmark incorporates three distinct audio feature sets as input to the classification models. Using raw audio signals directly is impractical due to their complexity and size, as they often contain redundant information. By performing feature extraction, we simplify the audio data, reduce dimensionality, improve training and inference efficiency, and minimise overfitting (Sharma et al., 2020; Tzanetakis, 2011). The feature sets used include acoustic features, mel frequency cepstral coefficients (MFCCs), and Wav2Vec2 feature vectors. Acoustic features and MFCCs were chosen for their widespread application in similar research (Paterson et al., 2024). Although Wav2Vec2 feature vectors have not previously been applied to throat cancer detection, they have shown effectiveness in identifying other pathologies (Wagner et al., 2023). To ensure consistency, all audio recordings were resampled to 16,000 Hz before feature extraction. This is the required sample rate for Wav2Vec2 and was applied across all feature sets.
Wav2Vec2 is a large speech model trained to perform automatic speech recognition (Baevski et al., 2020). In this work, we use Facebook’s XLSR-Wav2Vec2 model; this is a pre-trained model trained using CommonVoice and Multilingual LibriSpeech, two large multi-lingual datasets (Conneau et al., 2021; Ardila et al., 2020; Pratap et al., ). Feature vectors are extracted from the final convolutional layer of the Wav2Vec2 model (from here out referred to as FeatureStates). Since the feature embeddings have variable lengths corresponding to the varying durations of the audio recordings, we applied mean pooling to standardise the representation for each recording.
OpenSMILE is an open-source feature extraction toolkit (Eyben et al., 2010). In this study, we extract the eGeMAPSv02 feature set, which comprises 88 features, including low-level descriptors (frequency, amplitude, and spectral parameters), temporal features, and cepstral parameters (Eyben et al., 2016).
Mel frequency cepstral coefficients (MFCC) are derived from audio using fast Fourier transforms and Mel filtering. In this work, we extract 20 MFCC coefficients per recording. Since the length of the extracted MFCCs varies with the duration of the audio, we standardized the arrays by trimming those exceeding the average length and applying zero-padding to shorter ones. Finally, the standardized arrays were flattened before being input into the system.
3.2 Classification
Figure 3 shows the classification pipeline used in this work. Audio features and demographic/symptom data are preprocessed independently before being combined into a single input vector for the classification model.
For the audio, missing feature values are imputed as the mean of the respective feature. The audio features are then scaled, and feature selection is performed using a decision tree to manage the large number of audio features across all three feature sets. This process involves building a decision tree based on the audio features and then using the calculated feature importance to select only the most relevant features for input into the classifier.
[…]
For both the audio features and the demographic and symptom data, Z-score normalisation is performed to scale the features so that they have a mean of 0 and a standard deviation of 1.
Figure 3: The classification process used in this work. Where $\bar{x_{1}}$ if a vector of audio features and $\bar{x_{2}}$ is a vector of demographic/symptom data.
This work tests three commonly used classification algorithms: support vector machine (SVM), multilayered perceptron (MLP), and logistic regression. SVM and MLP were chosen due to their due to their established use in similar research areas (Degila et al., 2018; Miliaresi et al., 2021; Kim et al., 2024), while logistic regression was chosen for its simplicity. For each classifier, a grid search with 5-fold cross-validation was emploied during training to identify the optimal hyperparameters. The grid search explores various combinations of parameters defined in a parameter grid, training models for each combination. 5-fold cross-validation is used to evaluate the parameters by splitting the data into five subsets: the models are trained on four subsets and validated on the remaining one. This process is repeated, using a different validation subset each time. The parameter combination with the best average performance across all validation subsets is selected for the final model. Table 3 shows the parameter grids used for each algorithm.
Algorithm Hyperparamter Options
SVM C 0.1, 1, 10, 100, 1000
Gamma scale, auto, 1e-4, 1e-3, 0.01, 0.1, 1
Degree 2, 3, 4
Kernel linear, polynomial, rbf, sigmoid
MLP Hidden layer sizes (50,), (100,), (100, 50), (100, 100), (50, 50, 50)
Activation relu, tanh
Solver adam, sgd, lbfgs
Learning rate constant, invscaling, adaptive
Logistic Regression Penalty l1, l2, elasticnet, None
C 0.01, 0.1, 1, 10, 100
Solver newton-cg, lbfgs, liblinear, saga
Max iterations 100, 200, 300, 500
l1 ratio 0, 0.25, 0.5, 0.75, 1
Table 3: The parameter grid used in the cross-validation grid search for each of the algorithms.
To address the dataset’s imbalance of classes, class weighting was used in the SVM and logistic regression models (King and Zeng, 2001). The Synthetic Minority Oversampling Technique (SMOTE) was used for the MLP (Chawla et al., 2002). Balanced accuracy was used as the scoring metric for training across all three algorithms.
3.3 Evaluation Methods
Evaluating a model’s performance requires a multidimensional approach. While previous research in this field has primarily focused on predictive accuracy, it is crucial to consider additional metrics, especially for models intended for medical applications. This section discusses and suggests three aspects of model performance that should be assessed: predictive performance, fairness, and prediction time.
3.3.1 Predictive Performance
It’s important to choose contextually meaningful evaluation metrics for assessing classification models effectively. Previous works in this area have used a range of metrics to evaluate the performance of their models, with accuracy being the most commonly reported for binary classification. While accuracy is a very simple and well-understood metric, it is often unrepresentative of model performance on imbalanced datasets. To address this, we suggest using balanced accuracy (Equation 1), which averages the accuracy of each class, making it robust to class imbalance and providing a more comprehensive measure of overall of model performance. We also feel it is important to understand how a model classifies both the positive and negative cases. For this purpose, we suggest sensitivity (Equation 2) and specificity (Equation 3). These metrics provide insight into the model’s ability to correctly identify positive cases (sensitivity) and negative cases (specificity), offering a deeper understanding of its classification performance.
balanced $\text{balanced accuracy}=\frac{\sum_{i=0}^{n}\frac{TP_{i}}{TP_{i}+FN_{i}}}{n}$ (1)
$\text{sensitivity}=\frac{TP}{TP+FN}$ (2)
$\text{specificity}=\frac{TN}{TN+FP}$ (3)
In the given equations, TP stands for true positive, FN for false negative, TN for true negative, and FP for false positive. In Equation 1 $n$ is the number of classes, which in binary classification is two. Note that the positive class should be taken as malignant in these applications.
In addition to the above metrics, we also suggest the area under the receiver operating characteristic curve (AUROC). The ROC curve plots the true positive rate against the false positive rate at different decision thresholds. AUROC is particularly useful for evaluating models when the positive class is small compared to the negative class, which is common in the datasets used in this area.
3.3.2 Fairness
In medical application, it is important to assess the fairness of the algorithms being developed. As such we suggest conductng statistical tests to evaluate the relationship between classifier performance and patient demographics. The most commonly available patient demographics are gender and age of patients. For gender, we propose using a Fisher’s Exact Test to examine any association between classifier performance and patient gender (Sprent, 2011). We propose a Fisher’s Exact test as it is used for categorical variables and widely recommended for small sample sizes (Kim, 2016). For age, we suggest a t-test to compare performance across different age groups (Kalpić et al., 2011). Supplementary figures may also be helpful and should be included in supplementary material or code repositories. These figures should include more details on the correctly and incorrectly classified patients, such as stacked bar charts or boxplots.
3.3.3 Prediction Time
When applying these models in a healthcare setting, it is crucial to consider not only their accuracy but also their practicality. For a model to be effectively implemented, it must deliver predictions quickly to avoid delays in patient care. Therefore, reporting the inference times for individual audio files is essential. Models with longer inference times may be less suitable for clinical practice as they could delay patient referrals. These times should be reported end-to-end, including all preprocessing, feature extraction, and prediction.
4 Results and Discussion
All experiments were performed using Python 3.10.12 on a machine running Windows 11 with an Intel Core i7-1260P CPU with 16GB RAM. All code and models are available on github at https://github.com/mary-paterson/LaryngealCancerClassificationBenchmark
4.1 Predictive Performance
Table 4 shows the results of the classifiers on both the FEMH (holdout) and SVD (external) test sets. The bolded numbers indicate the best value for each metric on the test set. The results show that adding demographics consistently improves classifier performance. This improvement is expected, as the demographic differences between classes in both datasets align with patterns likely to occur in real-world applications.
On the holdout test set, when using voice only, FeatureStates achieved the best performance for both the SVM and MLP classifiers, while MFCC performs the best for logistic regression. On the external test set, FeatureStates outperformed all other features across all three algorithms. This outcome aligns with expectations, since Wav2Vec2, the source of FeatureStates, was trained on multiple datasets and designed to be generalisable across different recording environments and devices, making it best suited for both holdout and external test sets.
When demographics are used alongside voice for classification, OpenSmile performs best across all three algorithms for the holdout test set. On the external test set, OpenSmile performed best for MLP and logistic regression, while MFCC was best for SVM. It can be seen in Figure 4 that the audio features used had more of an effect on performance than the algorithm used with OpenSmile consistently outperforming FeatureStates and MFCC when combined with patient demographics.
The classifiers perform worse on the external test data than the holdout data. This is expected due to several differences between the datasets, including recording devices, recording environments, recording lengths, and demographics. The two datasets were also collected in different countries, so the speakers’ accents may have an effect, although this is limited by using the elongated vowel task. The benign groups also include different pathologies. It is difficult to disentangle these differences to understand what the leading cause of the performance reduction is or whether all of these factors have an effect. Figure 4 shows the difference in balanced accuracy between the holdout (FEMH) and external (SVD) test sets. Although the performance on the external dataset is generally worse, the confidence intervals overlap for most classifiers except for MFCC when using voice only for logistic regression.
Figure 4: The balanced accuracy for the holdout (FEMH) and external (SVD) test sets. 95% confidence intervals are shown; notably, the confidence interval cannot be seen for OpenSmile logistic regression with voice and demographics as input; this is because all samples were classified as malignant.
{adjustwidth}
-1.2cm-1.2cm FEMH (holdout test set) SVD (external test set) FeatureStates OpenSmile MFCC FeatureStates OpenSmile MFCC SVM Voice Balanced Accuracy 0.691 (0.593, 0.784) 0.680 (0.580, 0.772) 0.631 (0.531, 0.731) 0.628 (0.557, 0.696) 0.582 (0.497, 0.662) 0.532 (0.462, 0.613) Sensitivity 0.680 (0.480, 0.857) 0.640 (0.440, 0.818) 0.400 (0.206, 0.609) 0.763 (0.621, 0.900) 0.500 (0.341, 0.667) 0.289 (0.154, 0.455) Specificity 0.702 (0.665, 0.736) 0.720 (0.686, 0.754) 0.861 (0.835, 0.887) 0.493 (0.466, 0.521) 0.663 (0.637, 0.690) 0.775 (0.752, 0.799) AUC 0.760 (0.665, 0.845) 0.742 (0.676, 0.805) 0.701 (0.594, 0.807) 0.649 (0.562, 0.735) 0.615 (0.522, 0.711) 0.592 (0.503, 0.679) Voice + Demographics Balanced Accuracy 0.700 (0.594, 0.789) 0.752 (0.656, 0.837) 0.701 (0.599, 0.803) 0.609 (0.529, 0.691) 0.646 (0.566, 0.730) 0.681 (0.599, 0.759) Sensitivity 0.640 (0.429, 0.818) 0.760 (0.571, 0.923) 0.600 (0.400, 0.800) 0.526 (0.364, 0.688) 0.526 (0.370, 0.688) 0.605 (0.442, 0.756) Specificity 0.759 (0.726, 0.793) 0.743 (0.709, 0.776) 0.802 (0.770, 0.832) 0.691 (0.666, 0.717) 0.765 (0.741, 0.787) 0.758 (0.735, 0.781) AUC 0.773 (0.682, 0.855) 0.802 (0.743, 0.851) 0.807 (0.720, 0.881) 0.716 (0.642, 0.786) 0.772 (0.714, 0.829) 0.769 (0.699, 0.831) MLP Voice Balanced Accuracy 0.683 (0.576, 0.784) 0.632 (0.533, 0.727) 0.581 (0.478, 0.688) 0.628 (0.547, 0.705) 0.610 (0.529, 0.689) 0.548 (0.477, 0.616) Sensitivity 0.520 (0.318, 0.714) 0.560 (0.360, 0.750) 0.480 (0.280, 0.697) 0.526 (0.368, 0.676) 0.579 (0.417, 0.735) 0.763 (0.622, 0.889) Specificity 0.846 (0.817, 0.873) 0.704 (0.668, 0.740) 0.682 (0.646, 0.718) 0.729 (0.706, 0.753) 0.640 (0.615, 0.667) 0.333 (0.308, 0.358) AUC 0.754 (0.643, 0.857) 0.709 (0.637, 0.779) 0.638 (0.514, 0.746) 0.692 (0.607, 0.772) 0.633 (0.538, 0.727) 0.575 (0.489, 0.657) Voice + Demographics Balanced Accuracy 0.629 (0.522, 0.730) 0.755 (0.659, 0.841) 0.655 (0.556, 0.762) 0.669 (0.586, 0.748) 0.696 (0.612, 0.775) 0.657 (0.573, 0.742) Sensitivity 0.440 (0.227, 0.640) 0.760 (0.571, 0.923) 0.440 (0.250, 0.654) 0.553 (0.390, 0.704) 0.658 (0.488, 0.813) 0.500 (0.333, 0.667) Specificity 0.819 (0.786, 0.847) 0.750 (0.717, 0.784) 0.871 (0.844, 0.897) 0.786 (0.764, 0.808) 0.734 (0.708, 0.758) 0.815 (0.794, 0.835) AUC 0.758 (0.663, 0.833) 0.809 (0.753, 0.858) 0.805 (0.725, 0.881) 0.715 (0.627, 0.789) 0.758 (0.695, 0.822) 0.762 (0.695, 0.826) Logistic Regression Voice Balanced Accuracy 0.653 (0.551, 0.750) 0.663 (0.555, 0.761) 0.666 (0.563, 0.767) 0.619 (0.541, 0.689) 0.565 (0.487, 0.644) 0.528 (0.438, 0.603) Sensitivity 0.560 (0.360, 0.762) 0.640 (0.429, 0.833) 0.600 (0.400, 0.807) 0.711 (0.558, 0.849) 0.579 (0.428, 0.739) 0.526 (0.355, 0.676) Specificity 0.746 (0.713, 0.779) 0.685 (0.649, 0.721) 0.732 (0.698, 0.766) 0.528 (0.502, 0.557) 0.551 (0.524, 0.581) 0.530 (0.503, 0.555) AUC 0.724 (0.604, 0.834) 0.714 (0.637, 0.790) 0.719 (0.612, 0.825) 0.665 (0.578, 0.747) 0.603 (0.497, 0.704) 0.545 (0.439, 0.640) Voice + Demographics Balanced Accuracy 0.657 (0.544, 0.756) 0.797 (0.722, 0.859) 0.723 (0.623, 0.816) 0.661 (0.586, 0.735) 0.747 (0.683, 0.800) 0.701 (0.626, 0.774) Sensitivity 0.560 (0.345, 0.750) 0.880 (0.733, 1.000) 0.680 (0.500, 0.864) 0.711 (0.558, 0.849) 0.842 (0.714, 0.947) 0.684 (0.537, 0.829) Specificity 0.754 (0.722, 0.787) 0.715 (0.677, 0.749) 0.767 (0.732, 0.799) 0.612 (0.587, 0.639) 0.652 (0.627, 0.678) 0.718 (0.694, 0.743) AUC 0.733 (0.627, 0.827) 0.836 (0.767, 0.892) 0.813 (0.738, 0.883) 0.687 (0.607, 0.762) 0.777 (0.719, 0.828) 0.782 (0.721, 0.838)
Table 4: Classification results for each of the three acoustic features used alone and when combined with the demographics. These models are evaluated on both a holdout test set from the FEMH dataset and externally on the SVD dataset. Bold values are the highest for that metric and test set. SVM - support vector machine, MLP - multi-layered perceptron
The FEMH dataset includes data on 26 symptoms for each patient, which is not available in the SVD dataset. Consequently, classifiers incorporating symptom data cannot be evaluated on the external SVD dataset. Table 5 shows the results on the FEMH holdout test set when symptoms data are included as input to the classifier. Using this symptom data improves the classification performance across all algorithms and features. The best overall performance is found using logistic regression with symptom and demographic data combined with OpenSmile features, yielding a balanced accuracy of 83.7%.
We see an increase in accuracy when symptom data are included as input to the system. While this extra information is valuable, it is not easy to obtain. Although most clinics will obtain some medical history from their patients, this will vary between clinics, making it hard to generalise the system. There is also a concern that, should this be incorporated into an app for patient use, too many questions about their symptoms and medical history may be offputting and lead to inaccurate data entry as patients attempt to speed up the process.
Classifier Inputs Metric FEMH (holdout test dataset)
FeatureStates OpenSmile MFCC
SVM Voice + Demographics + Symptoms Balanced Accuracy 0.766 (0.671, 0.848) 0.807 (0.720, 0.872) 0.793 (0.707, 0.862)
Sensitivity 0.720 (0.526, 0.880) 0.840 (0.678, 0.962) 0.840 (0.667, 0.966)
Specificity 0.811 (0.780, 0.841) 0.775 (0.741, 0.807) 0.746 (0.714, 0.780)
AUROC 0.863 (0.799, 0.918) 0.893 (0.835, 0.937) 0.848 (0.796, 0.892)
Voice + Symptoms Balanced Accuracy 0.813 (0.720, 0.886) 0.814 (0.725, 0.891) 0.697 (0.597, 0.795)
Sensitivity 0.800 (0.609, 0.950) 0.800 (0.619, 0.950) 0.640 (0.438, 0.826)
Specificity 0.827 (0.796, 0.857) 0.828 (0.799, 0.856) 0.754 (0.722, 0.788)
AUROC 0.859 (0.785, 0.918) 0.866 (0.798, 0.918) 0.804 (0.723, 0.879)
MLP Voice + Demographics + Symptoms Balanced Accuracy 0.695 (0.596, 0.800) 0.776 (0.673, 0.859) 0.779 (0.676, 0.872)
Sensitivity 0.440 (0.241, 0.647) 0.720 (0.516, 0.885) 0.680 (0.483, 0.867)
Specificity 0.950 (0.931, 0.967) 0.831 (0.801, 0.859) 0.877 (0.852, 0.903)
AUROC 0.865 (0.787, 0.933) 0.892 (0.839, 0.937) 0.883 (0.828, 0.936)
Voice + Symptoms Balanced Accuracy 0.742 (0.641, 0.838) 0.792 (0.701, 0.873) 0.786 (0.698, 0.857)
Sensitivity 0.640 (0.440, 0.826) 0.760 (0.571, 0.920) 0.800 (0.619, 0.946)
Specificity 0.844 (0.817, 0.873) 0.824 (0.791, 0.851) 0.772 (0.739, 0.803)
AUROC 0.839 (0.764, 0.903) 0.889 (0.834, 0.936) 0.865 (0.802, 0.919)
Logistic Regression Voice + Demographics + Symptoms Balanced Accuracy 0.798 (0.711, 0.871) 0.837 (0.752, 0.901) 0.827 (0.754, 0.887)
Sensitivity 0.800 (0.633, 0.947) 0.840 (0.680, 0.966) 0.880 (0.737, 1.000)
Specificity 0.797 (0.765, 0.826) 0.833 (0.802, 0.862) 0.773 (0.741, 0.806)
AUROC 0.896 (0.844, 0.941) 0.918 (0.871, 0.954) 0.910 (0.863, 0.948)
Voice + Symptoms Balanced Accuracy 0.807 (0.723, 0.874) 0.786 (0.695, 0.869) 0.806 (0.719, 0.874)
Sensitivity 0.840 (0.677, 0.966) 0.760 (0.579, 0.923) 0.840 (0.667, 0.966)
Specificity 0.773 (0.739, 0.806) 0.813 (0.782, 0.842) 0.772 (0.738, 0.806)
AUROC 0.868 (0.797, 0.927) 0.879 (0.831, 0.922) 0.878 (0.821, 0.925)
Table 5: Classification results for the three acoustic features combined with the demographics and symptoms. These models could only be evaluated on a holdout test set from the FEMH dataset, as the external test set (SVD) does not include the required symptom data. Bold values are the highest for that metric and test set.
4.2 Fair Classification
In this section, we perform statistical tests comparing the correctly and incorrectly classified patients and their gender and age. First, a Fisher Exact test was used to compare the gender of correctly and incorrectly classified samples. This was done only on the FEMH holdout test set. Table 6 shows the p values from the Fisher Exact test for all classifiers. All of the p values are below the commonly used threshold of 0.05, meaning that there is a statistically significant relationship between the gender of the speaker and the performance of the classifier. This may be due to the imbalance of male and female participants across the malignant and benign groups. When investigating the patients that were correctly and incorrectly classified, we found male patients were misclassified more often than female patients across all models. This is likely because there are more male patients in both classes, whereas female patients are mainly in the benign class. Female patients with a malignant pathology were often misclassified as having a benign pathology due to this imbalance.
Model Input FeatureStates OpenSmile MFCC
SVM Voice 4.54e-33 2.44e-37 8.13e-22
Voice + Demographics 6.38e-40 1.17e-55 7.54e-40
Voice + Symptoms 1.48e-21 3.56e-22 1.36e-28
Voice + Demographics + Symptoms 5.72e-35 2.01e-15 1.06e-36
MLP Voice 2.60e-15 7.87e-52 1.10e-26
Voice + Demographics 3.26e-30 1.80e-56 7.56e-27
Voice + Symptoms 1.55e-15 1.08e-16 2.37e-19
Voice + Demographics + Symptoms 1.19e-11 2.79e-26 1.67e-21
Logistic Regression Voice 1.13e-22 1.50e-47 1.06e-13
Voice + Demographics 6.39e-25 6.72e-83 3.19e-56
Voice + Symptoms 3.13e-24 4.03e-18 1.45e-25
Voice + Demographics + Symptoms 4.51e-46 2.37e-37 5.52e-40
Table 6: The p values from the Fisher Exact test comparing the male and female patients and whether they were correctly or incorrectly classified. All models show significance at the 0.05 threshold.
We performed a t-test to compare the ages of the correctly and incorrectly classified patients. The p-values are shown in Table 7. For all models, the p-value is below the standard 0.05 threshold, indicating a significant difference between the ages of the correctly and incorrectly classified patients. Further investigation found that younger patients are more often correctly classified than older patients. This is likely because young patients are almost exclusively found in the benign group and are therefore easier to classify.
Model Input FeatureStates OpenSmile MFCC
SVM Voice 1.75e-11 1.39e-05 6.15e-04
Voice + Demographics 3.70e-23 2.33e-28 2.40e-29
Voice + Symptoms 9.17e-10 6.89e-06 1.52e-06
Voice + Demographics + Symptoms 1.34e-20 8.40e-20 9.08e-13
MLP Voice 2.19e-07 7.57e-05 1.23e-04
Voice + Demographics 4.84e-14 7.02e-27 7.61e-21
Voice + Symptoms 6.45e-06 9.45e-05 1.74e-04
Voice + Demographics + Symptoms 3.43e-09 3.86e-13 7.52e-11
Logistic Regression Voice 2.97e-12 7.09e-04 1.77e-02
Voice + Demographics 2.79e-18 8.65e-33 3.82e-28
Voice + Symptoms 7.77e-10 2.11e-10 4.94e-07
Voice + Demographics + Symptoms 3.95e-31 5.60e-21 7.10e-13
Table 7: The p values from the t-test comparing the age of the patients and whether they were correctly or incorrectly classified. All models show significance at the 0.05 threshold.
There is a statistical significance in the age and gender of correctly and incorrectly classified patients, even when voice alone is used as the system input. This suggests that there are audio features that may be indicative of a patient’s age and gender even when this information is not explicitly entered into the classifier. It is well known that the voices of males and females vary in pitch, although there can be a significant overlap (Latinus and Taylor, 2012). There is also an expected difference in voice as people age, with a change in pitch and the variation in pitch and amplitude (Rojas et al., 2020). These differences may cause some differences in classifier performance based on a patient’s sex or age without the information being specifically input into the system.
Graphs showing the difference in correctly and incorrectly classified male and female patients, as well as the age distribution of correctly and incorrectly classified patients, can be found in the GitHub repository and the appendix.
4.3 Prediction Time
In this section, we report the time taken to make a prediction for a single audio file for all of the presented models.
Figure 5 shows the time it takes to get a prediction from a single audio file, including feature extraction. The FeatureStates classifiers take significantly longer than the OpenSmile and MFCC classifiers due to the feature extraction. Overall, inference time does not vary significantly between the inputs or algorithms.
Figure 5: The time required to make a prediction from an audio file, including feature extraction. Outliers have been excluded from this figure.
Feature extraction takes the majority of the time taken to obtain a prediction. Figure 6 shows the inference times taken for samples after feature extraction. It is clear that even after feature extraction, the FeatureStates classifiers take the longest and MFCC the shortest amount of time, although the time taken for prediction is very short across all algorithms. The FeatureStates is the largest feature set, comprising 512 features. OpenSmile comprises 88, and MFCC comprises 20 coefficients but varies based on the recording length. When using the average length of an MFCC in the training set and flattening it, this comprises of 1540 features. The MFCCs are may be processed faster due to them being stored as numpy arrays rather than dataframes. However, since the difference is in the range of hundreds of seconds, this is unlikely to make a real-world difference.
We believe that since the total inference times, including feature extraction, are consistently less than 2 seconds, these inference times are short enough to be reasonably implemented into clinical practice.
Figure 6: The time required to make a prediction from an audio file after feature extraction. Outliers have been excluded from this figure.
5 Conclusion and Future Work
This work presents 36 models for classifying malignant and benign voice pathologies based on patients’ speech, demographics, and symptoms. We report the model’s performance, inference times, and statistical tests comparing classifier performance to patient demographics. All training and evaluation code, pre-trained models, and figure generation are available in a public repository for use as a baseline so that future work may be improved.
We find that while voice alone can be used to classify benign and malignant patients, the addition of demographic and symptom data improves classifier performance (69.1% balanced accuracy with voice only, 83.7% balanced accuracy with voice, demographics, and symptoms). We also find that while these classifiers may perform well on a holdout test set, performance degrades on external data.
In future work, we hope to improve these baseline models and their generalizability and robustness to improve their implementation in clinical practice. We hope to enhance the fairness of models in terms of patient’s age and gender. We hope that, as more datasets are release, these models may be further evaluated and enhanced.